OpenStudy (anonymous):

Hey I need help understanding this. What kind of verb would the word remained be?

OpenStudy (anonymous):

this is math lol

OpenStudy (anonymous):

And what is the difference between a transitive verb, an auxiliary verb?

OpenStudy (anonymous):

oops!

OpenStudy (anonymous):

can you please help me on kime then Iw ill help on your adn DRAW the answer please only in 2 grade

OpenStudy (anonymous):

What is your question

OpenStudy (anonymous):

ok what is 4.50*4.47 and draw it please so i can understadn it

OpenStudy (anonymous):

http://www.wolframalpha.com/ Just plug that question into there, and it will give you the graph. ( :

OpenStudy (anonymous):

ok let me see it very quickly

OpenStudy (anonymous):

I mean it will solve it

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=4.50*4.47+ here is the link

OpenStudy (anonymous):

And if you want to understand it, I don't hae time to help, but I know a great resource.

OpenStudy (anonymous):

ok it did not show draw it out for me can you? please

OpenStudy (anonymous):

4.50*4.47 4.50 *4.47 = 3250 + 1800 + 1800 = 6580. Now you must be thinking, wow! No way1 It can't be that big.. haha It isn't. Just go back the same number of decimal places so 6580 = 65.80 ( :

OpenStudy (anonymous):

Whoops! Sorry I got it wrong. One moment.

OpenStudy (anonymous):

The answer is 20.115

OpenStudy (anonymous):

I'm sorry it isn't easy to type it.

OpenStudy (anonymous):

its fine